How can I translate this to an equation? h$100 bills and t $20 bills total $400
I do not understand and then I have to graph this and use the graph to determine three different combinations of this.

Answer by mananth(16949)

h ----------$100 bills and
t----------- $20 bills
total $400
100h + 20t =400.............(1)
/20
5h+t=20
t= 20-5h
plot the points for various values of h
h=0, t=20 (0,20)
h=10,t=30( 4,0)

X axis hundreds
y axis tens
when x=4 y=0
when x=0, y=20
so you can find relation between x & y
m.ananth@hotmail.ca